# A Radio Map Approach for Reduced Pilot CSI Tracking in Massive MIMO Networks

### v3

@@ II-C The CSI Tracking Problem
Denote $\mathcal{Y}_{t}=(\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{\text{$t$}})$ as the sequence of sparse observations along the trajectory $\mathcal{P}_{t}=(\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{t})$ based on sensing matrices $\mathcal{A}_{t}=(\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{t})$, where the CSI sequence is denoted as $\mathcal{H}_{t}=({\bf h}_{1},{\bf h}_{2},\dots,{\bf h}_{t})$. The goal of this paper is to track the CSI $\mathcal{H}_{t}$ based on the sequence of sparse observations $\mathcal{Y}_{t}$. In particular, we propose to recover the trajectory $\mathcal{P}_{t}$ as an intermediate step to assist the CSI tracking via the radio map =$\mathcal{M}=\{(\mathbf{x},{\bf C}({\bf x}):\mathbf{x}\in\mathcal{X}\}$. In turn, we also need to construct or update the radio map $\mathcal{M}$ based on $\mathcal{Y}_{t}$ and $\mathcal{P}_{t}$.
@@ III-A Problem Reformulation
×$\times p({\bf h}_{1})\times\prod_{t=2}^{T}\mathbb{P}(\mathbf{p}_{\text{$t$}}|\mathbf{p}_{t-1})\mathbb{P}(\mathbf{p}_{1}).$ (6)
@@ III-B Radio-Map-Embedded Switching Kalman Filtering
Here, we develop an SKF approach to jointly track the positions and the CSI. We first find the distribution of the position $\mathbf{p}_{t}$ given the measurements up to time $t$ and the trajectory up to time $t-1$. Then, with the radio map $\mathcal{M}$ and the conditional distribution of $\mathbf{p}_{t}$, the CSI $\mathbf{h}_{t}$ is recursively tracked leveraging an SKF. Finally, the position $\mathbf{p}_{t}$ is estimated based on the maximum likelihood criterion.
@@ III-B1 Distribution of the Position
Let $\pi_{t}(\mathbf{p}_{t})\triangleq p(\mathbf{p}_{t}|\mathcal{Y}_{t},\mathcal{P}_{t-1})$ be the distribution of the position at time $t$ given the measurements up to time $t$ and the trajectory up to time $t-1$. Using Bayes’ theorem, for $t=2,3,\dots T$, $\pi_{t}(\mathbf{p}_{t})$ is derived as
@@
$=\frac{p(\mathbf{y}_{t}|\mathbf{p}_{t})\mathbb{P}(\mathbf{p}_{t}|\mathbf{p}_{t-1})}{\sum_{\mathbf{x}\in\mathcal{X}}p({\bf y}_{t}|\mathbf{x})\mathbb{P}(\mathbf{x}|\mathbf{p}_{t-1})},$ (10)
@@ III-B2 Radio-Map-Embedded CSI Tracking
First, we derive a CSI prediction and its corresponding error covariance at time $t$ based on the observations $\mathcal{Y}_{t-1}$ up to time $t-1$.
Define $\hat{\mathbf{h}}_{t-1}\triangleq\mathbb{E}\{\mathbf{h}_{t-1}|\mathcal{Y}_{t-1}\}$ as the Bayesian estimator of the CSI $\mathbf{h}_{t-1}$ based on the observations $\mathcal{Y}_{t-1}$ up to time $t-1$. Define $\hat{\mathbf{h}}_{t|t-1}\triangleq\mathbb{E}\{\mathbf{h}_{t}|\hat{\mathbf{h}}_{t-1}\}$ as a prediction of the CSI $\mathbf{h}_{t}$ based on $\hat{\mathbf{h}}_{t-1}$. Using the CSI dynamic model (4), we have
@@
Define $\mathbf{e}_{t|t-1}\triangleq\mathbf{h}_{t}-\hat{\mathbf{h}}_{t|t-1}$ as the channel prediction error of $\mathbf{h}_{t}$ based on the observations $\mathcal{Y}_{t-1}$ up to time $t-1$. Substituting (12) into $\mathbf{e}_{t|t-1}$, we can obtain
@@
$=\gamma\mathbf{e}_{t-1}+\sqrt{1-\gamma^{2}}{\bf u}_{t},$ (13)
@@
$=\gamma^{2}\mathbf{Q}_{t-1}+(1-\gamma^{2})\mathbb{E}\{\mathbf{u}_{t}\mathbf{u}_{t}^{\text{H}}|\mathbf{p}_{t-1}\},$ (14)
@@
$=\sum_{\mathbf{p}_{t}\in\mathcal{X}}\pi_{t}({\bf p}_{t})\times\mathbf{C}(\mathbf{p}_{t}),$ (15)
where recall from (4) that $\mathbf{u}_{t}|\mathbf{p}_{t}\sim\mathcal{CN}(0,\mathbf{C}(\mathbf{p}_{t}))$ implying $\mathbb{E}\{\mathbf{u}_{t}\mathbf{u}_{t}^{\mathrm{H}}|\mathbf{p}_{t}\}=\mathbf{C}(\mathbf{p}_{t})$, and $\pi_{t}({\bf p}_{t})$ captures the location distribution at time $t$ based on observations $\mathcal{Y}_{t}$ up to time $t$ and the user trajectory $\mathcal{P}_{t-1}$ up to time $t-1$. It is observed that the radio map $\mathcal{M}$ plays a critical role in constructing the distribution of the hidden position variable $\mathbf{p}_{t}$ and the error covariance $\mathbf{Q}_{t|t-1}$ for the channel prediction.
@@
$\hat{\mathbf{h}}_{t}$ $=\hat{\mathbf{h}}_{t|t-1}+\mathbf{K}_{t}(\mathbf{y}_{t}-\mathbf{A}_{t}\hat{\mathbf{h}}_{t|t-1}),$ (16)
@@ III-B3 User Position Tracking
$=\frac{p(\mathbf{h}_{t}|\mathbf{h}_{t-1},\mathbf{p}_{t})\mathbb{P}(\mathbf{p}_{t}|\mathbf{p}_{t-1})}{\sum_{\mathbf{x}\in\mathcal{X}}p(\mathbf{h}_{t}|\mathbf{h}_{t-1},\mathbf{x})\mathbb{P}(\mathbf{x}|\mathbf{p}_{t-1})},$ (21)
@@
$=\mathop{\mbox{argmax}}\limits_{\mathbf{p}_{t}\in\mathcal{X}}\frac{\exp(-\frac{1}{2}\boldsymbol{\Delta}_{t}^{\mathrm{H}}((1-\gamma^{2})\mathbf{C}(\mathbf{p}_{t}))^{-1}\boldsymbol{\Delta}_{t})}{\sqrt{(2\pi(1-\gamma^{2}))^{N_{\mathrm{t}}}|\mathbf{C}(\mathbf{p}_{t})|}}\mathbb{P}(\mathbf{p}_{t}|\mathbf{p}_{t-1}),$ (22)
@@ III-C1 Differential Entropy
From (18), the error covariance $\mathbb{E}\{(\mathbf{h}_{t}-\hat{\mathbf{h}}_{t})(\mathbf{h}_{t}-\hat{\mathbf{h}}_{t})^{\text{H}}\}$ of $\hat{\mathbf{h}}_{t}$ given the measurements $\mathcal{Y}_{t}$ up to time $t$ is captured in $\mathbf{Q}_{t}$, which implies that $\hat{\mathbf{h}}_{t}\sim\mathcal{CN}(\mathbf{h}_{t},\mathbf{Q}_{t}).$ Following this, the differential entropy of $\hat{\mathbf{h}}_{t}$ is derived as [27, Theorem 8.4.1]
@@ III-C2 Solution for the Adaptive Sensing Matrix $\mathbf{A}_{t}$
$=|\mathbf{Q}_{t|t-1}^{-1}||\mathbf{I}+\sigma_{\text{n}}^{-2}\mathbf{A}_{t}\mathbf{Q}_{t|t-1}^{-1}\mathbf{A}_{t}^{\text{H}}|,$ (29)
@@
${\bf A}_{t}=[{\bf w}_{t-1,1},\mathbf{w}_{t-1,2},\dots,\mathbf{w}_{t-1,m},\dots,\mathbf{w}_{t-1,M}]^{\mathrm{H}},$ (32)
@@
The adaptive sensing matrices can be incorporated into the proposed radio-map-embedded SKF framework for efficient channel sensing. However, the design of $\mathbf{A}_{t}$ relies on the error covariance $\mathbf{Q}_{t|t-1}$, which is based on the position distribution $\pi_{t}(\mathbf{p}_{t})$ calculated using the measurement $\mathbf{y}_{t}$. To address this implementation issue, the transition probability $\mathbb{P}(\mathbf{p}_{t}|\mathbf{p}_{t-1})$ is treated as the prior position distribution before observing $\mathbf{y}_{t}$. Hence, the error covariance $\mathbf{Q}_{t|t-1}$ is first computed using $\mathbb{P}(\mathbf{p}_{t}|\mathbf{p}_{t-1})$, and then refined via $\pi_{t}(\mathbf{p}_{t})$ as detailed in Step 2) of Algorithm 1.
@@ IV-B1 Unbiased Estimator of the Covariance
$\boldsymbol{\varphi}_{t}=\mathbf{A}_{t}^{\mathrm{H}}\mathbf{y}_{t}=\mathbf{A}_{t}^{\mathrm{H}}\mathbf{A}_{t}\mathbf{h}_{t}+\mathbf{A}_{t}^{\text{H}}\mathbf{n}_{t}=\boldsymbol{\Phi}_{t}\mathbf{h}_{t}+\mathbf{A}_{t}^{\text{H}}\mathbf{n}_{t},$ (35)
@@
[…]
$-\frac{\sigma_{\mathrm{n}}^{2}(M\bar{N}_{\mathrm{t}}\boldsymbol{\Omega}_{\mathrm{A}}(\mathbf{x}_{i})-N_{\mathrm{t}}(N_{\mathrm{t}}-M)\mathrm{tr}(\boldsymbol{\Omega}_{\mathrm{A}}(\mathbf{x}_{i}))\mathbf{I})}{M(N_{\mathrm{t}}M+N_{\mathrm{t}}-2)},$ (37)
@@
[…]
@@ IV-B2 Upper Bound of Reconstruction Error
[…]
@@
$\|\hat{\mathbf{C}}^{\prime}(\mathbf{x}_{i})-\mathbf{C}(\mathbf{x}_{i})\|_{2}$ $\leq\frac{\kappa}{\sqrt{|\mathcal{T}_{i}|}}\|\mathbf{C}(\mathbf{x}_{i})\|_{2}\left(S_{1}+S_{2}+S_{3}\right),$ (39)
@@
Proposition 5 provides key insights into the factors influencing the reconstruction error of sample covariance. First, if $|\mathcal{T}_{i}|$ is large relative to $N_{\mathrm{t}}^{2}/M^{2}$ and the logarithmic factors are ignored, the leading terms in the error bound (39) is $\tilde{\mathcal{O}}(\|\mathbf{C}(\mathbf{x}_{i})\|_{2}(\sqrt{N_{\mathrm{t}}R^{2}/(|\mathcal{T}_{i}|M^{2})}+N_{\mathrm{t}}R/(|\mathcal{T}_{i}|M))),$ where the notation $\tilde{\mathcal{O}}(n)$ represents the asymptotic growth rate of a function $n$ while suppressing dependencies on logarithmic factors. Thus, increasing the number of observation samples reduces error terms inversely, making it crucial to accumulate sufficient samples. Second, the upper bound in (39) decreases inversely with $M$, and thus, increasing $M$ may result in a reduction of the reconstruction error. Additionally, the rank of the true channel covariance highlights the difficulty of estimating channel covariance in NLOS environments with rich multipath effects. By contrast, in LOS environments, where $R$ is small, reconstruction accuracy improves naturally, requiring fewer pilot signals (see Table. IV).

### v5

@@ II-C The CSI Tracking Problem
Denote $\mathcal{Y}_{t}=(\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{\text{$t$}})$ as the sequence of sparse observations along the trajectory $\mathcal{P}_{t}=(\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{t})$ based on sensing matrices $\mathcal{A}_{t}=(\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{t})$, where the CSI sequence is denoted as $\mathcal{H}_{t}=({\bf h}_{1},{\bf h}_{2},\dots,{\bf h}_{t})$. The goal of this paper is to track the CSI $\mathcal{H}_{t}$ based on the sequence of sparse observations $\mathcal{Y}_{t}$. In particular, we propose to recover the trajectory $\mathcal{P}_{t}$ as an intermediate step to assist the CSI tracking via the radio map $\mathcal{M}=\{(\mathbf{x},{\bf C}({\bf x}):\mathbf{x}\in\mathcal{X}\}$. In turn, we also need to construct or update the radio map $\mathcal{M}$ based on $\mathcal{Y}_{t}$ and $\mathcal{P}_{t}$.
@@ III-A Problem Reformulation
$\times p({\bf h}_{1})\times\prod_{t=2}^{T}\mathbb{P}(\mathbf{p}_{\text{$t$}}|\mathbf{p}_{t-1})\mathbb{P}(\mathbf{p}_{1}).$ (6)
@@ III-B Radio-Map-Embedded Switching Kalman Filtering
Here, we develop an SKF approach to jointly track the positions and the CSI. We first find the distribution of the position $\mathbf{p}_{t}$ given the measurements up to time $t$ and the trajectory up to time t1t-1. Then, with the radio map $\mathcal{M}$ and the conditional distribution of $\mathbf{p}_{t}$, the CSI $\mathbf{h}_{t}$ is recursively tracked leveraging an SKF. Finally, the position $\mathbf{p}_{t}$ is estimated based on the maximum likelihood criterion.
@@ III-B1 Distribution of the Position
Let $\pi_{t}(\mathbf{p}_{t})\triangleq p(\mathbf{p}_{t}|\mathcal{Y}_{t},\mathcal{P}_{t-1})$ be the distribution of the position at time $t$ given the measurements up to time $t$ and the trajectory up to time t1t-1. Using Bayes’ theorem, for $t=2,3,\dots T$, $\pi_{t}(\mathbf{p}_{t})$ is derived as
@@
$=\frac{p(\mathbf{y}_{t}|\mathbf{p}_{t})\mathbb{P}(\mathbf{p}_{t}|\mathbf{p}_{t-1})}{\sum_{\mathbf{x}\in\mathcal{X}}p({\bf y}_{t}|\mathbf{x})\mathbb{P}(\mathbf{x}|\mathbf{p}_{t-1})}$ (10)
@@ III-B2 Radio-Map-Embedded CSI Tracking
First, we derive a CSI prediction and its corresponding error covariance at time $t$ based on the observations $\mathcal{Y}_{t-1}$ up to time t1t-1.
Define $\hat{\mathbf{h}}_{t-1}\triangleq\mathbb{E}\{\mathbf{h}_{t-1}|\mathcal{Y}_{t-1}\}$ as the Bayesian estimator of the CSI $\mathbf{h}_{t-1}$ based on the observations $\mathcal{Y}_{t-1}$ up to time t1t-1. Define $\hat{\mathbf{h}}_{t|t-1}\triangleq\mathbb{E}\{\mathbf{h}_{t}|\hat{\mathbf{h}}_{t-1}\}$ as a prediction of the CSI $\mathbf{h}_{t}$ based on $\hat{\mathbf{h}}_{t-1}$. Using the CSI dynamic model (4), we have
@@
Define $\mathbf{e}_{t|t-1}\triangleq\mathbf{h}_{t}-\hat{\mathbf{h}}_{t|t-1}$ as the channel prediction error of $\mathbf{h}_{t}$ based on the observations $\mathcal{Y}_{t-1}$ up to time t1t-1. Substituting (12) into $\mathbf{e}_{t|t-1}$, we can obtain
@@
$=\gamma\mathbf{e}_{t-1}+\sqrt{1-\gamma^{2}}{\bf u}_{t}$ (13)
@@
$=\gamma^{2}\mathbf{Q}_{t-1}+(1-\gamma^{2})\mathbb{E}\{\mathbf{u}_{t}\mathbf{u}_{t}^{\text{H}}|\mathbf{p}_{t-1}\}$ (14)
@@
$=\sum_{\mathbf{p}_{t}\in\mathcal{X}}\pi_{t}({\bf p}_{t})\times\mathbf{C}(\mathbf{p}_{t})$ (15)
where recall from (4) that $\mathbf{u}_{t}|\mathbf{p}_{t}\sim\mathcal{CN}(0,\mathbf{C}(\mathbf{p}_{t}))$ implying $\mathbb{E}\{\mathbf{u}_{t}\mathbf{u}_{t}^{\mathrm{H}}|\mathbf{p}_{t}\}=\mathbf{C}(\mathbf{p}_{t})$, and $\pi_{t}({\bf p}_{t})$ captures the location distribution at time $t$ based on observations $\mathcal{Y}_{t}$ up to time $t$ and the user trajectory $\mathcal{P}_{t-1}$ up to time t1t-1. It is observed that the radio map $\mathcal{M}$ plays a critical role in constructing the distribution of the hidden position variable $\mathbf{p}_{t}$ and the error covariance $\mathbf{Q}_{t|t-1}$ for the channel prediction.
@@
$\hat{\mathbf{h}}_{t}$ $=\hat{\mathbf{h}}_{t|t-1}+\mathbf{K}_{t}(\mathbf{y}_{t}-\mathbf{A}_{t}\hat{\mathbf{h}}_{t|t-1})$ (16)
@@ III-B3 User Position Tracking
$=\frac{p(\mathbf{h}_{t}|\mathbf{h}_{t-1},\mathbf{p}_{t})\mathbb{P}(\mathbf{p}_{t}|\mathbf{p}_{t-1})}{\sum_{\mathbf{x}\in\mathcal{X}}p(\mathbf{h}_{t}|\mathbf{h}_{t-1},\mathbf{x})\mathbb{P}(\mathbf{x}|\mathbf{p}_{t-1})}$ (21)
@@
$=\mathop{\mbox{argmax}}\limits_{\mathbf{p}_{t}\in\mathcal{X}}\frac{\exp(-\frac{1}{2}\boldsymbol{\Delta}_{t}^{\mathrm{H}}((1-\gamma^{2})\mathbf{C}(\mathbf{p}_{t}))^{-1}\boldsymbol{\Delta}_{t})}{\sqrt{(2\pi(1-\gamma^{2}))^{N_{\mathrm{t}}}|\mathbf{C}(\mathbf{p}_{t})|}}\mathbb{P}(\mathbf{p}_{t}|\mathbf{p}_{t-1})$ (22)
@@ III-C1 Differential Entropy
From (18), the error covariance $\mathbb{E}\{(\mathbf{h}_{t}-\hat{\mathbf{h}}_{t})(\mathbf{h}_{t}-\hat{\mathbf{h}}_{t})^{\text{H}}\}$ of $\hat{\mathbf{h}}_{t}$ given the measurements $\mathcal{Y}_{t}$ up to time $t$ is captured in $\mathbf{Q}_{t}$, which implies that $\hat{\mathbf{h}}_{t}\sim\mathcal{CN}(\mathbf{h}_{t},\mathbf{Q}_{t}).$ Following this, the differential entropy of $\hat{\mathbf{h}}_{t}$ is derived as [28, Theorem 8.4.1]
@@ III-C2 Solution for the Adaptive Sensing Matrix $\mathbf{A}_{t}$
$=|\mathbf{Q}_{t|t-1}^{-1}||\mathbf{I}+\sigma_{\text{n}}^{-2}\mathbf{A}_{t}\mathbf{Q}_{t|t-1}^{-1}\mathbf{A}_{t}^{\text{H}}|$ (29)
@@
${\bf A}_{t}=[{\bf w}_{t-1,1},\mathbf{w}_{t-1,2},\dots,\mathbf{w}_{t-1,m},\dots,\mathbf{w}_{t-1,M}]^{\mathrm{H}}$ (32)
@@
The adaptive sensing matrices can be incorporated into the proposed radio-map-embedded SKF framework for efficient channel sensing. However, the design of $\mathbf{A}_{t}$ relies on the error covariance $\mathbf{Q}_{t|t-1}$, which is based on the position distribution $\pi_{t}(\mathbf{p}_{t})$ calculated using the measurement $\mathbf{y}_{t}$. To address this implementation issue, the transition probability $\mathbb{P}(\mathbf{p}_{t}|\mathbf{p}_{t-1})$ is treated as the prior position distribution before observing $\mathbf{y}_{t}$. Hence, the error covariance $\mathbf{Q}_{t|t-1}$ is first computed using $\mathbb{P}(\mathbf{p}_{t}|\mathbf{p}_{t-1})$, and then refined via $\pi_{t}(\mathbf{p}_{t})$ as detailed in Step 2a and Step 2c of Algorithm 1, respectively.
@@ IV-B1 Unbiased Estimator of the Covariance
$\boldsymbol{\varphi}_{t}=\mathbf{A}_{t}^{\mathrm{H}}\mathbf{y}_{t}=\mathbf{A}_{t}^{\mathrm{H}}\mathbf{A}_{t}\mathbf{h}_{t}+\mathbf{A}_{t}^{\text{H}}\mathbf{n}_{t}=\boldsymbol{\Phi}_{t}\mathbf{h}_{t}+\mathbf{A}_{t}^{\text{H}}\mathbf{n}_{t}$ (35)
@@
$\text{$\hat{\mathbf{C}}$}(\mathbf{x}_{i})=$ $\frac{M(\bar{N}_{\mathrm{t}}\hat{\boldsymbol{\Omega}}_{\text{y}}(\mathbf{x}_{i})-(N_{\mathrm{t}}-M)\mathrm{tr}(\hat{\boldsymbol{\Omega}}_{\text{y}}(\mathbf{x}_{i}))\mathbf{I})}{N_{\mathrm{t}}(N_{\mathrm{t}}M+N_{\mathrm{t}}-2)}$
$-\frac{\sigma_{\mathrm{n}}^{2}(M\bar{N}_{\mathrm{t}}\boldsymbol{\Omega}_{\mathrm{A}}(\mathbf{x}_{i})-N_{\mathrm{t}}(N_{\mathrm{t}}-M)\mathrm{tr}(\boldsymbol{\Omega}_{\mathrm{A}}(\mathbf{x}_{i}))\mathbf{I})}{M(N_{\mathrm{t}}M+N_{\mathrm{t}}-2)}$ (37)
@@
The unbiased estimator $\text{$\hat{\mathbf{C}}$}(\mathbf{x}_{i})$ in Proposition 4 is composed of a scaled version of the sample covariance of the CSI measurements, with an additional correction term that accounts for the noise covariance. Consider the special case where $M=N_{\text{t}}$ and $\boldsymbol{\Phi}_{t}=\mathbf{A}_{t}^{\text{H}}\mathbf{A}_{t}=\mathbf{I}$, the estimator simplifies to $\hat{\mathbf{C}}(\mathbf{x}_{i})=\sum_{t\in\mathcal{T}_{i}}(\boldsymbol{\varphi}_{t}\boldsymbol{\varphi}_{t}^{\mathrm{H}}-\sigma_{\mathrm{n}}^{2}\mathbf{I})/|\mathcal{T}_{i}|$. Next, we can take the expectation of $\text{$\hat{\mathbf{C}}$}(\mathbf{x}_{i})$ and obtain that $\mathbb{E}\{\hat{\mathbf{C}}(\mathbf{x}_{i})\}=\mathbb{E}\{\sum_{t\in\mathcal{T}_{i}}((\mathbf{h}_{t}+\mathbf{A}_{t}^{\text{H}}\mathbf{n}_{t})(\mathbf{h}_{t}+\mathbf{A}_{t}^{\text{H}}\mathbf{n}_{t})^{\mathrm{H}}-\sigma_{\mathrm{n}}^{2}\mathbf{I})/|\mathcal{T}_{i}|\}=\mathbf{C}(\mathbf{x}_{i})$.
@@ IV-B2 Upper Bound of Reconstruction Error
Consider the high SNR case, where $\sigma_{\text{n}}=0$, and the observed CSI is simplified as $\mathbf{y}_{t}=\mathbf{A}_{t}\mathbf{h}_{t}$. Based on (37), the unbiased estimator $\text{$\hat{\mathbf{C}}$}(\mathbf{x}_{i})$ in (37) simplifies to
@@
$\|\hat{\mathbf{C}}^{\prime}(\mathbf{x}_{i})-\mathbf{C}(\mathbf{x}_{i})\|_{2}$ $\leq\frac{\kappa}{\sqrt{|\mathcal{T}_{i}|}}\|\mathbf{C}(\mathbf{x}_{i})\|_{2}\left(S_{1}+S_{2}+S_{3}\right)$ (39)
@@
Proposition 5 provides key insights into the factors influencing the reconstruction error of sample covariance. First, if $|\mathcal{T}_{i}|$ is large relative to $N_{\mathrm{t}}^{2}/M^{2}$ and the logarithmic factors are ignored, the leading terms in the error bound (39) is $\tilde{\mathcal{O}}(\|\mathbf{C}(\mathbf{x}_{i})\|_{2}(\sqrt{N_{\mathrm{t}}R^{2}/(|\mathcal{T}_{i}|M^{2})}+N_{\mathrm{t}}R/(|\mathcal{T}_{i}|M))),$ where the notation $\tilde{\mathcal{O}}(n)$ represents the asymptotic growth rate of a function $n$ while suppressing dependencies on logarithmic factors. Thus, increasing the number of observation samples reduces error terms inversely, making it crucial to accumulate sufficient samples. Second, the upper bound in (39) decreases inversely with $M$, and thus, increasing $M$ may result in a reduction of the reconstruction error. Additionally, the rank of the true channel covariance highlights the difficulty of estimating channel covariance in NLOS environments with rich multipath effects. By contrast, in LOS environments, where $R$ is small, reconstruction accuracy improves naturally, requiring fewer pilot observations (see Table IV).